OpenStudy (anonymous):

For the function f(x) = 2 – 3x, find f(4). A. –10 B. 12 C. 14 D. –5 A?

OpenStudy (anonymous):

just sub in 4 for x and solve

OpenStudy (anonymous):

f(4) = 2 - 3(4) f(4) = 2 - 12 f(4) = -10 you are correct...it is A

OpenStudy (anonymous):

It's A , =2-(3•4) =-10

OpenStudy (anonymous):

Okay thank you I just wanted to double check! (: